OpenStudy (anonymous):

PLEASE HELP! the ordered pairs (1,2) (2,4) (3,8) (4,16) and (5,32) represent a function. what is a rule that represents this function?

OpenStudy (anonymous):

looks like you are taking the number \(2\) to the power of the first coordinate. i.e. \[2^1=2,2^2=4,2^3=8,2^4=16\] and so on

OpenStudy (anonymous):

my choices are Y=x^2 y=2x y=2^x y= x +2

OpenStudy (anonymous):

@satellite73

OpenStudy (emilylauren):

it would be C then

OpenStudy (anonymous):

Thanks so much ,could u you help with one more?

OpenStudy (anonymous):

@EmilyLauren

OpenStudy (anonymous):

1. (1,6) (2,12) (3,18) (4,24) (5,30) Y=6x Y= (x+6)^2 y=x^6 y=6^x

OpenStudy (ranga):

If you look at each ordered pair you can see that the y coordinate is six times the x coordinate.

OpenStudy (emilylauren):

yes, one second

OpenStudy (emilylauren):

the answer is A 6x1=6 6x2=12 6x3=18 etc.